OpenStudy (tennistar):

Using complete sentences, compare and contrast the ways to factor the "sum of two squares" and the "difference of two squares." Provide an explanation and example of each, including any similarities and differences.

OpenStudy (anonymous):

The sum of squares is in the form of (a ± b)^2 = a^2 ± 2ab + b^2 To factor it, just take the square root of the first term and the third term and use the sign of the middle term to keep them apart. THe difference in squares is in the form of (a + b)(a - b) = a^2 - b^2 TO factor it, just have two binomials with a plus in one and a minus in the other (in between a and b)